What is the pH of a solution prepared by adding 0.2g of Na2CO3 and 0.2g of NaHCO3 to water and bringing to a volume of 1liter? (pKa2=10.25)


1
Expert's answer
2021-04-19T08:50:56-0400

Moles of "Na_2CO_3=" "0.2g\over 105.99g/mol" "=0.00199moles"

Molarity="0.00199moles\over 1L" "U=1.90\times 10^{-3} M"

Moles of "NaHCO_3=" "0.2g\over 84.007g/mol" "0.00239moles"

Molarity of "=" "0.00239moles\over 1L" "=2.39\times 10^{-3} M"

"pH=pKa_2+log[HCO_3^{2-}\/pCO-2]"

"pH=10.25+log2.39\times 10^{-3}"

"pH=10.25+(-0.924)=9.32"